OpenStudy (anonymous):

A country's population in 1994 was 195 million. In 2002 it was 199 million. Estimate the population in 2016 using exponential growth.

OpenStudy (campbell_st):

you need a model like \[A = P \times e^{kt}\] A = the future population, P = initial population, k = growth constant t = time in years so at t = 0 the population P = 195 and t= 8, the population is 199 this can be used to find k, the growth constant \[199 = 195 \times e^{8k}\] solve for k... when you get k use the formula \[A = 195 \times e^{kt}\] using the k value you found and t = 12, which is the time between 1994 and 2012 hope it helps